# DVLO: Deep Visual-LiDAR Odometry with Local-to-Global Feature Fusion and Bi-Directional Structure Alignment

### v1

@@ Abstract
Information inside visual and LiDAR data is well complementary derived from the fine-grained texture of images and massive geometric information in point clouds. However, it remains challenging to explore effective visual-LiDAR fusion, mainly due to the intrinsic data structure inconsistency between two modalities: Images are regular and dense, but LiDAR points are unordered and sparse. To address the problem, we propose a local-to-global fusion network with bi-directional structure alignment. To obtain locally fused features, we project points onto image plane as cluster centers and cluster image pixels around each center. Image pixels are pre-organized as pseudo points for image-to-point structure alignment. Then, we convert points to pseudo images by cylindrical projection (point-to-image structure alignment) and perform adaptive global feature fusion between point features with local fused features. Our method achieves state-of-the-art performance on KITTI odometry and FlyingThings3D scene flow datasets compared to both single-modal and multi-modal methods. Codes will be released upon publication.
@@ 1 Introduction
Visual/LiDAR odometry is a fundamental task in the field of robotics and computer vision, which aims to estimate the relative pose transformation between two consecutive images or point clouds. It is widely applied in autonomous driving [60, 47], SLAM [48, 58], and virtual/augmented reality [7, 21], etc. Recently, the multi-modal odometry [11, 51, 62] has been gaining increasing attention because it can take advantage of complementary information from different modalities. In this paper, we mainly focus on the investigation of the visual-LiDAR fusion odometry.
Previous visual-LiDAR odometry works can be classified into two categories: traditional methods [51, 55, 12, 11, 5] and learning-based methods [45, 62, 42]. Traditional methods accomplish the odometry task through a pipeline consisting of feature extraction, frame-to-frame feature matching, motion estimation, and optimization [5, 55]. However, these methods suffer from inaccurate pose estimation because of the poor quality and low resolution of extracted features [5, 51]. With the development of deep learning, some methods [45, 13] attempt to utilize CNN-based methods for visual-LiDAR fusion and pose estimation. However, the receptive field of feature fusion is limited by the stride and kernel size of CNN. To enlarge the receptive field for multi-modal fusion, attention-based methods are recently proposed, which leverage cross-attention mechanism for the multi-modal fusion [62, 42]. Attention-based methods can fuse multi-modal features globally and establish the cross-frame association with larger receptive fields because of their long-range dependencies. However, due to the quadratic computational complexity, attention-based methods commonly require larger computational consumption and longer inference time [29], which challenges the real-time applications [42]. Moreover, previous learning-based methods mostly adopt the only feature-level fusion strategy as illustrated in Fig. 1 a), which fails to capture fine-grained pixel-to-point correspondences [25]. Recently, some networks [23, 62] design point-to-image projection and local feature aggregation as in Fig. 1 b). However, their performances are still limited by the intrinsic data structure misalignment between sparse LiDAR points and dense camera pixels [25].
@@
To address these problems, we propose a novel local-to-global fusion network with bi-directional structure alignment in Fig. 1 c). Our fusion module consists of two parts: 1) An image is first viewed as a set of pseudo points inspired by [30] for fine-grained local fusion with LiDAR points (image-to-point structure alignment). 2) Point clouds are also converted into pseudo images by cylindrical projection for global adaptive fusion (point-to-image structure alignment). Specifically, a novel clustering-based local fusion module (Local Fuser) is designed to perform local fine-grained feature fusion. We first project LiDAR points onto the image plane based on their coordinate calibration matrices to find the corresponding pixels as cluster centers. Meanwhile, for the image-to-point structure alignment, image pixels are reshaped as a set of pseudo points [30]. Then, within a certain scope of each cluster center, we aggregate pseudo point features to generate the local fused features based on the similarities to the cluster centers dynamically. In global fusion module, we project point clouds onto the cylindrical surface to obtain pseudo images. Then, an adaptive fusion mechanism is leveraged to merge the above local fused image features and point (pseudo image) features for global fusion. It is worth noting that our fusion module is hierarchically utilized at multi-scale feature maps between images and points. The local fusion module can provide more fine-grained point-to-pixel correspondence information, while the global fusion has larger receptive fields and achieves more global information interaction.
@@
    Extensive experiments on the KITTI odometry dataset [10, 9] demonstrate that our method outperforms all recent deep LiDAR, visual, and visual-LiDAR fusion odometry works on most sequences. Furthermore, our fusion strategy can generalize well to other multi-modal tasks, like scene flow estimation, even surpassing recent SOTA method CamLiRAFT.
@@ 2 Related Work
Deep Visual Odometry.   Recently, learning-based methods have shown impressive performance in the visual odometry field [49]. The pioneering work [17] uses deep neural networks for odometry estimation with the prediction of both speed and direction for individual images. PoseNet [15] initially employs Convolution Neural Networks (CNNs) to extract features from the input image and then estimate the pose. Based on PoseNet, Szegedy et al. [43] replace the linear regression and softmax layers with a fully connected layer. Additionally, they utilize the reprojection error as a loss function to enhance the pose supervision. DeepVO [50] applies deep recurrent neural networks to capture the temporal dynamics and interdependency information of sequences, thereby facilitating the estimation of ego-motion. TartanVO [53] enhances the generalization by incorporating cosine similarity loss and normalized distance loss. Li et al. [20] utilize the knowledge distillation technique based on pre-trained visual-LiDAR odometry as a teacher for guiding the training of the visual odometry. NeRF-VO [32] improves the geometric accuracy of the scene representation by optimizing a set of keyframe poses and the underlying dense geometry through training the radiance field with volume rendering.
@@
Deep LiDAR Odometry.   In contrast to visual odometry, deep LiDAR odometry is still a challenging task because of the large number, irregularity, and sparsity of raw LiDAR points [46]. Nicolai et al. [33] first introduce the deep learning technique into LiDAR odometry. They project 3D LiDAR points onto the 2D plane to obtain 2D depth images, and then employ 2D learning methods for pose estimation. DeepPCO [52] projects point clouds to panoramic depth images and applies two sub-networks to estimate the translation and rotation respectively. LO-Net [22] also converts points to 2D format by projection and uses the normal of each 3D point and dynamical masks to further improve the performance. PWCLO [47] introduces the PWC [41] structure for the LiDAR odometry task, which hierarchically refines estimated poses by an iterative warp-refinement module. EfficientLO [46] proposes a projection-aware operator for improving the efficiency of LiDAR odometry. TransLO [26] designs a window-based masked point transformer to enhance global feature embeddings and remove outliers. DELO [1] introduces partial optimal transportation of LiDAR descriptors and predictive uncertainty for robust pose estimation. NeRF-LOAM [8] applies a neural radiation field to the LiDAR odometry system, showing excellent generalization capabilities across various environments.
@@
Viusal-LiDAR Odometry.   Recently, there has been an increasing focus on visual-LiDAR odometry, which takes advantage of both 2D texture and 3D geometric features. Existing visual-LiDAR odometry can be classified into two categories: traditional methods and learning-based methods. For traditional methods, DEMO [57] employs LiDAR points to provide depth information for the RGB images. V-LOAM [59] leverages the high-frequency estimated poses from visual odometry as a motion prior for low-frequency LiDAR odometry, resulting in refined motion estimation. LIMO [11] utilizes depth information derived from LiDAR points to alleviate the scale uncertainty that is intrinsic to monocular visual odometry. PL-LOAM [12] provides a pure visual motion tracking method and a novel scale correction algorithm. DV-LOAM [51] is a SLAM framework including a two-stage direct visual odometry module, a LiDAR mapping module with considerations of dynamic objects, and a parallel global and local search loop closure detection module. SDV-LOAM [55] combines a semi-direct visual odometry with an adaptive sweep-to-map LiDAR odometry to tackle the challenges of 3D-2D depth correlation. For learning-based methods, MVL-SLAM [3] employs the RCNN network architecture, fusing RGB images and multi-channel depth images from 3D LiDAR points. LIP-Loc [39] proposes a pre-training strategy for cross-modal localization, which utilizes contrastive learning to jointly train image and point encoders. However, the problem of natural data structure inconsistency between points and images is not fully considered before. To the best of our knowledge, our work is the first visual-LiDAR odometry network with bi-directional structure alignment.
@@ 3.1 Overall Architecture
The overall architecture of our proposed DVLO is illustrated in Fig. 2. Given two point clouds $PC_{S}$, $PC_{T}\in\mathbb{R}^{N\times 3}$ and their corresponding monocular camera images $I_{S}$, $I_{T}\in\mathbb{R}^{H\times W\times 3}$ from a pair of consecutive frames, the goal of our odometry is to estimate the relative pose including both rotation quarternion $q\in\mathbb{R}^{4}$ and translation vector $t\in\mathbb{R}^{3}$ between two frames.
In detail, we first feed the camera images and LiDAR points into the hierarchical feature extraction module in Sec. 3.2 to obtain multi-level image features and point features. Then, in Sec. 3.3 and Sec. 3.4, we delve into the details of our designed local-to-global fusion module. Finally, poses are estimated and iteratively refined from fused features in Sec. 3.5.
Figure 2: The pipeline of our proposed DVLO. We propse a novel Local-to-Global (LoGo) fusion module, which consists of a clustering-based Local Fuser and an adaptive Global Fuser. The pose is initially regressed from the cost volume of the coarsest fused features and then refined iteratively from fused features in shallower layers.
@@ 3.2 Hierarchical Feature Extraction
where $x$, $y$, $z$ are the raw 3D coordinates of point cloud. $u$, $v$ are the corresponding 2D pixel positions on the projected pseudo image. $\Delta\theta$ and $\Delta\phi$ are horizontal and vertical resolutions of the LiDAR sensor, respectively. To make the best use of geometric information of raw 3D points, we fill each projected 2D position with its corresponding raw $x$, $y$, $z$ 3D coordinates. In this case, LiDAR points can not only be converted into the pseudo image structure [22] for better alignment and global feature fusion with images in Sec. 3.4, but also retain the original 3D geometric information for effective feature extraction. Then, pseudo images of size $H_{P}\times W_{P}\times 3$ in Fig. 2 will be fed into the hierarchical feature extraction module [46] to extract multi-level point features $F_{P}\in\mathbb{R}^{H_{P}\times W_{P}\times D}$, where $D$ is the number of channels of the pseudo image features.
@@ 3.3 Local Fuser Module
Figure 3: Our designed Local-to-Global (LoGo) Fusion module. We project points onto the image plane based on the coordinate system transformation matrix as cluster centers and convert image into set of pseudo points. Then, we locally aggregate pseudo point features based on the similarities to each cluster center.
Pseudo Point Clustering.   We first project LiDAR points onto the image plane to obtain their corresponding 2D coordinates $x^{\prime}$, $y^{\prime}$ in the image coordinate system as cluster centers. The center feature $F_{c}\in\mathbb{R}^{N\times C}$ is computed by the bilinear interpolation on $F_{I}$ based on $x^{\prime}$, $y^{\prime}$. Then, we divide all pseudo points into several clusters according to the pair-wise cosine similarities between center features $F_{c}$ and pseudo point features $F_{pp}$. Here, we allocate each pseudo point to the most similar center, resulting in $N$ clusters. For efficiency, following Swin Transformer [29], we use the region partition while computing similarity.
@@
where $F_{pp}^{j}$ is the j-th pseudo point’s feature. $s_{ij}$ is the similarity score between the $j$-th pseudo point and $i$-th cluster center. $\alpha$ and $\beta$ are learnable scalars to scale and shift the similarity. $sigmoid\left(\cdot\right)$ is a sigmoid function to re-scale the similarity to (0, 1). $X$ is the normalization factor. Since we project LiDAR points onto the image plane as cluster centers and aggregate features for each center, local fused feature $F_{L}\in\mathbb{R}^{N\times C}$ has the same dimension with original LiDAR points. Therefore, we can also reshape the local fused feature $F_{L}$ like a pseudo image with size of $H_{P}\times W_{P}\times C$ as the input of the Global Fuser module.
@@ 3.4 Global Fuser Module
Since local feature fusion is only conducted within a partitioned region, above Local Fuser module has a limited receptive field. To expand the receptive field for sufficient feature fusion, we introduce the global adaptive fusion mechanism [35] between local fused feature $F_{L}$ and point (pseudo image) feature $F_{P}$ as shown in Fig. 3. Our Gloabl Fuser can enable the global information interaction between two modalities. The global modeling ability facilitates to recognize dynamics and occlusion, because they introduce inconsistent global motions [27].
@@ 3.6 Loss Function
Table 1: Comparison with different odometry networks on the KITTI odometry dataset [9]. $t_{rel}$ and $r_{rel}$ mean the average sequence translational RMSE (%) and the average sequence rotational RMSE (/100m) respectively on 00-10 subsequences in the length of 100, 200, …, 800m. Best results for each sequence are bold, and the second best results are underlined. represents the model is trained on the 00-08 sequences.
@@ 4.1 KITTI Odometry Dataset
We evaluate our DVLO on the KITTI odometry dataset [9], which is a widely used benchmark for the evaluation of odometry and SLAM system. The dataset consists of 22 sequences of LiDAR point clouds and their corresponding stereo images. In this paper, we only use the monocular left camera image for the fusion with LiDAR sensor. Since the ground truth pose (trajectory) is only available for sequences 00-10, we utilize these sequences for training and testing.

### v3

@@ Abstract
Information inside visual and LiDAR data is well complementary derived from the fine-grained texture of images and massive geometric information in point clouds. However, it remains challenging to explore effective visual-LiDAR fusion, mainly due to the intrinsic data structure inconsistency between two modalities: Image pixels are regular and dense, but LiDAR points are unordered and sparse. To address the problem, we propose a local-to-global fusion network (DVLO) with bi-directional structure alignment. To obtain locally fused features, we project points onto the image plane as cluster centers and cluster image pixels around each center. Image pixels are pre-organized as pseudo points for image-to-point structure alignment. Then, we convert points to pseudo images by cylindrical projection (point-to-image structure alignment) and perform adaptive global feature fusion between point features and local fused features. Our method achieves state-of-the-art performance on KITTI odometry and FlyingThings3D scene flow datasets compared to both single-modal and multi-modal methods. Codes are released at https://github.com/IRMVLab/DVLO.
@@ 1 Introduction
Visual/LiDAR odometry is a fundamental task in the field of computer vision and robotics, which estimates the relative pose transformation between two consecutive images or point clouds. It is widely applied in autonomous driving [19, 55, 71, 62], SLAM [8, 72, 9, 61], navigation [33, 35, 34] etc. Recently, the multi-modal odometry [15, 57, 73] has gained increasing attention because it can take advantage of complementary information from different modalities and possess the strong robustness for asymmetric sensor degradation [10].
Previous visual-LiDAR odometry works can be classified into two categories: traditional methods [57, 64, 16, 15] and learning-based methods [53, 73, 51]. Traditional methods accomplish the odometry task through a pipeline consisting of feature extraction, frame-to-frame feature matching, motion estimation, and optimization [64]. However, these methods suffer from inaccurate pose estimation because of the poor quality and low resolution of extracted features [57]. With the development of deep learning, some methods [53, 17] attempt to utilize CNN-based methods for visual-LiDAR fusion and pose estimation. However, the receptive field of feature fusion is limited by the stride and kernel size of CNN. To enlarge the receptive field for multi-modal fusion, attention-based methods are recently proposed, which leverage the cross-attention mechanism for the multi-modal fusion [73, 51]. Attention-based methods can fuse multi-modal features globally and establish the cross-frame association with larger receptive fields because of their long-range dependencies. However, due to the quadratic computational complexity, attention-based methods commonly require larger computational consumption and longer inference time [37], which challenges the real-time applications [51]. Moreover, previous learning-based methods mostly adopt the only feature-level fusion strategy as illustrated in Fig. 1 a), which fails to capture fine-grained pixel-to-point correspondences [29]. Recently, some networks [27, 73] design point-to-image projection and local feature aggregation as in Fig. 1 b). However, their performances are still limited by the intrinsic data structure misalignment between sparse LiDAR points and dense camera pixels [29].
@@
To address these problems, we propose a novel local-to-global fusion network (DVLO) with bi-directional structure alignment in Fig. 1 c). Our fusion module consists of two parts: 1) An image is first viewed as a set of pseudo points inspired by [38] for fine-grained local fusion with LiDAR points (image-to-point structure alignment). 2) Point clouds are also converted into pseudo images by cylindrical projection for global adaptive fusion (point-to-image structure alignment). Specifically, a novel clustering-based local fusion module (Local Fuser) is designed to perform local fine-grained feature fusion. We first project LiDAR points onto the image plane based on their coordinate calibration matrices to find the corresponding pixels as cluster centers. Meanwhile, for the image-to-point structure alignment, image pixels are reshaped as a set of pseudo points [38]. Then, within a certain scope of each cluster center, we aggregate pseudo point features to generate the local fused features based on the similarities to the cluster centers dynamically. In the global fusion module, we project point clouds onto the cylindrical surface to obtain pseudo images. Then, an adaptive fusion mechanism is leveraged to merge the above local fused image features and point (pseudo image) features for global fusion. It is worth noting that our fusion module is hierarchically utilized at multi-scale feature maps between images and points. The local fusion module can provide more fine-grained point-to-pixel correspondence information, while the global fusion has a larger receptive field and achieves more global information interaction.
@@
    Extensive experiments on the KITTI odometry dataset [14, 13] demonstrate that our method outperforms all recent deep LiDAR, visual, and visual-LiDAR fusion odometry works on most sequences. Furthermore, our fusion strategy can generalize well to other multi-modal tasks, like scene flow estimation, even surpassing recent SOTA method CamLiRAFT [29].
@@ 2 Related Work
Deep Visual Odometry. Recently, learning-based methods have shown impressive performance in the visual odometry field [22, 21, 56, 40, 45]. The pioneering work [22] uses deep neural networks for odometry estimation with the prediction of both speed and direction for individual images. PoseNet [21] initially employs Convolution Neural Networks (CNNs) to extract features from the input image and then estimate the pose. DeepVO [56] applies deep recurrent neural networks to capture the temporal dynamics and interdependency information of sequences, thereby facilitating the estimation of ego-motion. Li et al. [25] utilize the knowledge distillation technique based on pre-trained visual-LiDAR odometry as a teacher for guiding the training of the visual odometry. Deng et al. [10] propose a long-term visual SLAM system with map prediction and dynamics removal. NeRF-VO [40] improves the geometric accuracy of the scene representation by optimizing a set of keyframe poses and the underlying dense geometry through training the radiance field with volume rendering [59, 60].
@@
Deep LiDAR Odometry. In contrast to visual odometry, deep LiDAR odometry is still a challenging task because of the large number, irregularity, and sparsity of raw LiDAR points [54, 5, 69]. Nicolai et al. [41] first introduce the deep learning technique into LiDAR odometry. They project 3D LiDAR points onto the 2D plane to obtain 2D depth images, and then employ 2D learning methods for pose estimation. DeepPCO [58] projects point clouds to panoramic depth images and applies two sub-networks to estimate the translation and rotation respectively. LO-Net [26] also converts points to 2D format by projection and uses the normal of each 3D point and dynamical masks to further improve the performance. PWCLO [55] introduces the PWC [50] structure for the LiDAR odometry task, which hierarchically refines estimated poses by an iterative warp-refinement module. EfficientLO [54] proposes a projection-aware operator for improving the efficiency of LiDAR odometry. TransLO [30] designs a window-based masked point transformer to enhance global feature embeddings and remove outliers. DELO [1] introduces partial optimal transportation of LiDAR descriptors and predictive uncertainty for robust pose estimation. NeRF-LOAM [7] applies a neural radiation field to the LiDAR odometry system, showing excellent generalization capabilities across various environments.
@@
Viusal-LiDAR Odometry. Recently, there has been an increasing focus on visual-LiDAR odometry, which takes advantage of both 2D texture and 3D geometric features. Existing visual-LiDAR odometry can be classified into two categories: traditional methods and learning-based methods. For traditional methods, V-LOAM [67] leverages the high-frequency estimated poses from visual odometry as a motion prior for the low-frequency LiDAR odometry, resulting in refined motion estimation. LIMO [15] utilizes depth information derived from LiDAR points to alleviate the scale uncertainty that is intrinsic to monocular visual odometry. PL-LOAM [16] provides a pure visual motion tracking method and a novel scale correction algorithm. DV-LOAM [57] is a SLAM framework including a two-stage direct visual odometry module, a LiDAR mapping module with considerations of dynamic objects, and a parallel global and local search loop closure detection module. SDV-LOAM [64] combines the semi-direct visual odometry with an adaptive sweep-to-map LiDAR odometry to tackle the challenges of 3D-2D depth correlation. For learning-based methods, MVL-SLAM [3] employs the RCNN network architecture, fusing RGB images and multi-channel depth images from 3D LiDAR points. LIP-Loc [48] proposes a pre-training strategy for cross-modal localization, which utilizes contrastive learning to jointly train image and point encoders. However, the problem of natural data structure inconsistency between points and images has not been fully considered before. To the best of our knowledge, our work is the first visual-LiDAR odometry network with bi-directional structure alignment.
@@ 3.1 Overall Architecture
The overall architecture of our proposed DVLO is illustrated in Fig. 2. Given two point clouds $PC_{S}$, $PC_{T}\in\mathbb{R}^{N\times 3}$ and their corresponding monocular camera images $I_{S}$, $I_{T}\in\mathbb{R}^{H\times W\times 3}$ from a pair of consecutive frames, the goal of our odometry is to estimate the relative pose including both rotation quaternion $q\in\mathbb{R}^{4}$ and translation vector $t\in\mathbb{R}^{3}$ between two frames.
In detail, we first feed the camera images and LiDAR points into the hierarchical feature extraction module in Sec. 3.2 to obtain multi-level image features and point features. Then, in Sec. 3.3 and Sec. 3.4, we delve into the details of our designed local-to-global fusion module. Finally, poses are estimated and iteratively refined from fused features as in Sec. 3.5.
Figure 2: The pipeline of our proposed DVLO. We propose a novel Local-to-Global (LoGo) fusion module, which consists of a clustering-based Local Fuser and an adaptive Global Fuser. The pose is initially regressed from the cost volume of the coarsest fused features and then refined iteratively from fused features in shallower layers.
@@ 3.2 Hierarchical Feature Extraction
where $x$, $y$, $z$ are the raw 3D coordinates of point cloud. $u$, $v$ are the corresponding 2D pixel positions on the projected pseudo image. $\Delta\theta$ and $\Delta\phi$ are horizontal and vertical resolutions of the LiDAR sensor, respectively. To make the best use of geometric information of raw 3D points, we fill each projected 2D position with its corresponding original 3D coordinates. In this case, LiDAR points can not only be converted into the pseudo image structure [26] for better alignment and global feature fusion with images in Sec. 3.4, but also retain the original 3D geometric information for effective feature extraction. Then, pseudo images of size $H_{P}\times W_{P}\times 3$ in Fig. 2 will be fed into the hierarchical feature extraction module [54] to extract multi-level point features $F_{P}\in\mathbb{R}^{H_{P}\times W_{P}\times D}$, where $D$ is the number of channels of the pseudo image features.
@@ 3.3 Local Fuser Module
Figure 3: Our designed Local-to-Global (LoGo) Fusion module. We project points onto the image plane based on the coordinate system transformation matrix as cluster centers and convert the image into a set of pseudo points. Then, we locally aggregate pseudo point features based on the similarities to each cluster center.
Pseudo Point Clustering. We first project LiDAR points onto the image plane to obtain their corresponding 2D coordinates $x^{\prime}$ and $y^{\prime}$ in the image coordinate system as cluster centers. The center feature $F_{c}\in\mathbb{R}^{N\times C}$ is computed by the bilinear interpolation on $F_{I}$ based on $x^{\prime}$, $y^{\prime}$. Then, we divide all pseudo points into several clusters according to the pair-wise cosine similarities between center features $F_{c}$ and pseudo point features $F_{pp}$. Here, we allocate each pseudo point to the most similar center, resulting in $N$ clusters. For efficiency, following Swin Transformer [37], we use the region partition while computing similarity.
@@
where $F_{pp}^{j}$ is the j-th pseudo point’s feature. $s_{ij}$ is the similarity score between the $j$-th pseudo point and $i$-th cluster center. $\alpha$ and $\beta$ are learnable scalars to scale and shift the similarity. $sigmoid\left(\cdot\right)$ is a sigmoid function to re-scale the similarity to (0, 1). $X$ is the normalization factor. Since we project LiDAR points onto the image plane as cluster centers and aggregate features for each center, local fused feature $F_{L}\in\mathbb{R}^{N\times C}$ has the same dimension with original LiDAR points. Therefore, we can also reshape the local fused feature $F_{L}$ like a pseudo image with the size of $H_{P}\times W_{P}\times C$ as the input of the Global Fuser module.
@@ 3.4 Global Fuser Module
Since local feature fusion is only conducted within a partitioned region, the above Local Fuser module has a limited receptive field. To expand the receptive field for sufficient feature fusion, we introduce the global adaptive fusion mechanism [43] between local fused feature $F_{L}$ and point (pseudo image) feature $F_{P}$ as shown in Fig. 3. Our Global Fuser can enable the global information interaction between two modalities, which facilitates to recognize dynamics and occlusion because they introduce inconsistent global motions [31].
@@ 3.6 Loss Function
Table 1: Comparison with different odometry networks on the KITTI odometry dataset [13]. $t_{rel}$ and $r_{rel}$ mean the average sequence translational RMSE (%) and the average sequence rotational RMSE (/100m) respectively on 00-10 subsequences in the length of 100, 200, …, 800m. The best results are bold, and the second best results are underlined. represents the model is trained on the 00-08 sequences.
@@ 4.1 KITTI Odometry Dataset
We evaluate our DVLO on the KITTI odometry dataset [13], which is a widely used benchmark for the evaluation of odometry and SLAM system. The dataset consists of 22 sequences of LiDAR point clouds and their corresponding stereo images. In this paper, we only use the monocular left camera image for the fusion with the LiDAR sensor. Since the ground truth pose (trajectory) is only available for sequences 00-10, we utilize these sequences for training and testing.